If (a,2) is a point of the graph of y=3x+5, what is a?
Answer by harpazo(655)

We know that y is 2.
y = 3x + 5
2 = 3a + 5
2-5 = 3a
-3 = 3a
-3/3 = a
-1 = a
Done.